# Quantitative local inversion and a minimax depth barrier for canonical systems from fixed-height Weyl-Schur seam data

### v1

@@ 1.1. Setup
We often write traceless real-symmetric perturbations using a complex parameter $q=a+ib\in\mathbb{C}\simeq\mathbb{R}^{2}$ via
@@ 2.1. Tail extension and Weyl coefficient
Write $\Phi_{H}(\Lambda,0;z)=\begin{pmatrix}A&B\ C&D\end{pmatrix}$. On the tail $\Lambda,\infty)$ the system is $Y^{\prime}=-\frac{z}{2}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}Y$. For $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ the HH–square-integrable solution is proportional to $e^{iz(s-\Lambda)/2}u_{+}$, where $u_{+}=(1,i)^{\top}$. Thus the Weyl solution satisfies $Y(\Lambda)\parallel u_{+}$.
@@ 2.2. A perturbation formula for the Weyl function as a spectral pairing
Differentiate the identity $\Phi_{\varepsilon}g_{\varepsilon}=u_{+}$ at $\varepsilon=0$ to obtain $\dot{g}=-\Phi^{-1}\dot{\Phi}\,g$ (dot = $\partial_{\varepsilon}|_{\varepsilon=0}$). The standard variation-of-constants formula gives
@@ 3. Base linearization at the free Hamiltonian
Let $H_{0}\equiv\tfrac{1}{2}I$ on $0,\Lambda]$ (and tail). Define the vectors
@@ 3.4. Quadratic remainder bounds in $L^{1}$
Step 2: an exact Duhamel identity. Let $\Phi_{\Delta}(s,t):=\Phi_{H_{0}+\Delta H}(s,t;z)$ and $\Phi_{0}(s,t):=\Phi_{H_{0}}(s,t;z)$. A standard variation-of-constants argument (equivalently, subtract the integral equations (23) for $H_{0}+\Delta H$ and $H_{0}$ and solve for $\Phi_{\Delta}-\Phi_{0}$) gives, for $0\leq t\leq s\leq\Lambda$,
@@
The first term is quadratic by (32). For the second term, note that $\left|\delta n-i\delta d\right|\lesssim|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}}$ by the Lipschitz bound $\|\Phi_{\Delta}-\Phi_{0}\|_{\mathrm{op}}\leq|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}}$ and the definitions of $n_{\Delta},d_{\Delta}$, while $\left|d_{0}-d_{\Delta}\right|\lesssim\|\Delta H\|_{L^{1}}$ and $\left|d_{\Delta}\right|^{-1}\lesssim e^{-(\operatorname{Im}z)\Lambda/2}$ by (33). Combining these bounds with (32) and inflating constants yields (30) with the stated $C(z,\Lambda)$. Finally, the Cayley transform $w\mapsto(w-i)/(w+i)$ has uniformly bounded first and second derivatives on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ (e.g. by direct differentiation and $|w+i|\geq 1$ for $w\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$), so the same quadratic bound transfers from $m$ to $v$, giving (31). ∎
@@ 4.2. A well-posed design-minimax bound (uniform in real nodes)
In particular, if $s_{\star}$ is the maximal depth start among the basis functions, then $\Sigma^{*}_{M}\lesssim\sqrt{M}\,e^{-\eta s_{\star}}$ with an explicit constant.
@@ 7.2. First and second variations and explicit bounds
$C_{1}(z,\Lambda):=|z|\Lambda\,e^{2|z|\Lambda},\qquad C_{2}(z,\Lambda):=2|z|^{2}\Lambda^{2}\,e^{3|z|\Lambda}.$
@@
$|z|^{2}\int_{0}^{\Lambda}\!\!\int_{0}^{s}e^{|z|(\Lambda-s)}\|\Delta H_{1}(s)\|_{\mathrm{op}}e^{|z|(s-u)}\|\Delta H_{2}(u)\|_{\mathrm{op}}e^{|z|u}\,du\,ds\leq|z|^{2}\Lambda^{2}e^{|z|\Lambda}\,\|\Delta\theta_{1}\|_{2}\|\Delta\theta_{2}\|_{2},$
and we again inflate constants to the stated $C_{2}$. ∎
@@ 10. A base-point-free exponential barrier from spectral smoothing
This is the de Branges correspondence in the trace-normed class: the map $H\mapsto m_{H}$ is a bijection between trace-normed half-line Hamiltonians and Herglotz–Nevanlinna functions, with uniqueness up to equality a.e.; see [10, Thm. 2.2]. ∎
@@
which is Herglotz on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ but has logarithmic branch points at $\pm 1$, hence cannot be meromorphic on $\mathbb{C}$. Therefore imposing a free tail after $\Lambda$ cuts down the realizable class strictly.
@@ 10.3. Herglotz representation and Poisson smoothing
i.e. $\operatorname{Im}m(\cdot+i\eta)$ is the Poisson integral (up to $c_{2}\eta$) of the spectral measure.
@@
are Herglotz–Nevanlinna. Hence each $m_{\pm}$ is the Weyl coefficient of a (unique trace-normed) canonical system on $0,\infty)$.
@@ 10.5. A quantitative minimax obstruction for de-smoothing
For $\eta>0$, define the Poisson kernel
@@ 11. An operator-theoretic corollary: compact perturbations of the free Dirac system
acting in $L^{2}_{H}$. This is the standard Dirac-type interpretation of canonical systems; the Weyl function $m_{H,\Lambda}$ is exactly the Weyl–Titchmarsh function of $\tau_{H}$ for the boundary condition at $s=0$ and the $L^{2}_{H}$ condition at $\infty$. (For background, see de Branges’ theory and modern accounts such as [10, 20].)
@@
$f(x):=\frac{1}{\pi}\operatorname{Im}m(x+i0),\qquad g_{\eta}(x):=\operatorname{Im}m(x+i\eta)=(\mathcal{P}_{\eta}f)(x),$
where the last identity is the Herglotz/Poisson representation of Theorem˜10.8. Then the map
@@
The Poisson identity $g_{\eta}=\mathcal{P}_{\eta}f$ is Theorem˜10.8. The two-point construction and deterministic minimax lower bound are exactly Theorem˜10.12. Finally, Theorem˜10.1 realizes the constructed Herglotz functions as Weyl functions of trace-normed canonical systems. ∎

### v2

# arXiv:2601.06128v2[math.GM] 01 Mar 2026

@@ 1.1. Setup
We often write real-symmetric traceless perturbations using a complex parameter $q=a+ib\in\mathbb{C}\simeq\mathbb{R}^{2}$ via
@@ 2.1. Tail extension and Weyl coefficient
Write $\Phi_{H}(\Lambda,0;z)=\begin{pmatrix}A&B\ C&D\end{pmatrix}$. On the tail $\Lambda,\infty)$ the system is $Y^{\prime}=-\frac{z}{2}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}Y$. For $z\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ the HH–square-integrable solution is proportional to $e^{iz(s-\Lambda)/2}u_{+}$, where $u_{+}=(1,i)^{\top}$. Thus the Weyl solution satisfies $Y(\Lambda)\parallel u_{+}$ (this is exactly the square-integrable tail mode for $\operatorname{Im}z>0$).
@@ 2.2. A perturbation formula for the Weyl function as a spectral pairing
Differentiate the identity $\Phi_{\varepsilon}g_{\varepsilon}=u_{+}$ at $\varepsilon=0$ to obtain $\dot{g}=-\Phi^{-1}\dot{\Phi}\,g$ (dot = $\partial_{\varepsilon}|_{\varepsilon=0}$). The standard variation-of-constants formula gives; see [5].; see [5].
@@ 3. Base linearization at the free Hamiltonian
Let $H_{0}\equiv\tfrac{1}{2}I$ on $0,\Lambda]$ (and tail). Define the eigenvectors
@@ 3.4. Quadratic remainder bounds in $L^{1}$
Step 2: an exact Duhamel identity. Let $\Phi_{\Delta}(s,t):=\Phi_{H_{0}+\Delta H}(s,t;z)$ and $\Phi_{0}(s,t):=\Phi_{H_{0}}(s,t;z)$. A standard variation-of-constants argument (equivalently, subtract the integral equations (23) for $H_{0}+\Delta H$ and $H_{0}$ and see [5]) gives, for $0\leq t\leq s\leq\Lambda$, solve for $\Phi_{\Delta}-\Phi_{0}$) gives, for $0\leq t\leq s\leq\Lambda$,
@@
The first term is quadratic by (32). For the second term, note that $\left|\delta n-i\delta d\right|\leq 4|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}}$ by the Lipschitz bound $\|\Phi_{\Delta}-\Phi_{0}\|_{\mathrm{op}}\leq|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}}$ and the definitions of $n_{\Delta},d_{\Delta}$, while $\left|d_{0}-d_{\Delta}\right|\leq 2|z|e^{2|z|\Lambda}\|\Delta H\|_{L^{1}}$ and $\left|d_{\Delta}\right|^{-1}\leq\tfrac{4}{3}e^{-(\operatorname{Im}z)\Lambda/2}$ by (33). Combining these bounds with (32) and inflating constants yields (30) with the stated $C(z,\Lambda)$. Finally, the Cayley transform $w\mapsto(w-i)/(w+i)$ has uniformly bounded first and second derivatives on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ (e.g. by direct differentiation and $|w+i|\geq 1$ for $w\in\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$), so the same quadratic bound transfers from $m$ to $v$, giving (31). ∎
@@ 4.2. A well-posed design-minimax bound (uniform in real nodes)
In particular, if $s_{\star}$ is the maximal depth start among the basis functions, then
$\Sigma^{*}_{M}\leq\sqrt{M}\,e^{-\eta s_{\star}}\Bigl(\left|\varphi_{j_{\star}}(s_{\star})\right|+\left|\varphi_{j_{\star}}(\Lambda)\right|+\|\varphi_{j_{\star}}^{\prime}\|_{L^{1}(s_{\star},\Lambda)}\Bigr).$
@@ 7.2. First and second variations and explicit bounds
$C_{1}(z,\Lambda):=|z|\Lambda\,e^{2|z|\Lambda},\qquad C_{2}(z,\Lambda):=|z|^{2}\Lambda^{2}\,e^{3|z|\Lambda}.$
@@
$|z|^{2}\int_{0}^{\Lambda}\!\!\int_{0}^{s}e^{|z|(\Lambda-s)}\|\Delta H_{1}(s)\|_{\mathrm{op}}e^{|z|(s-u)}\|\Delta H_{2}(u)\|_{\mathrm{op}}e^{|z|u}\,du\,ds\leq\frac{1}{2}|z|^{2}\Lambda^{2}e^{|z|\Lambda}\,\|\Delta\theta_{1}\|_{2}\|\Delta\theta_{2}\|_{2},$
Since there are two such symmetric terms, their sum is bounded by $|z|^{2}\Lambda^{2}e^{|z|\Lambda}\|\Delta\theta_{1}\|_{2}\|\Delta\theta_{2}\|_{2}$, and we again inflate the exponential constant to the stated $C_{2}$. ∎
@@ 10. A base-point-free exponential barrier from spectral smoothing
This is the de Branges correspondence in the trace-normed class: the map $H\mapsto m_{H}$ is a bijection between trace-normed half-line Hamiltonians and Herglotz–Nevanlinna functions, with uniqueness up to equality a.e.; see [9, Thm. 2.2]. ∎
@@
which is Herglotz on $\{z\in\mathbb{C}:\ \operatorname{Im}z>0\}$ but has logarithmic branch points at $\pm 1$ (for the principal branch with a cut along [1,1][-1,1]), hence is not meromorphic on $\mathbb{C}$. Therefore imposing a free tail after $\Lambda$ cuts down the realizable class strictly.
@@ 10.3. Herglotz representation and Poisson smoothing
i.e. $\operatorname{Im}m(\cdot+i\eta)$ is the Poisson integral (up to $c_{2}\eta$) of the spectral measure. Equivalently, with the normalized Poisson kernel $P_{\eta}(u):=\frac{1}{\pi}\frac{\eta}{u^{2}+\eta^{2}}$, one has
$\frac{1}{\pi}\Big(\operatorname{Im}m_{H,\Lambda}(x+i\eta)-c_{2}\eta\Big)=\int_{\mathbb{R}}P_{\eta}(x-t)\,d\rho(t).$
In particular, if $c_{2}=0$ and $d\rho(t)=f(t)\,dt$, then $\frac{1}{\pi}\operatorname{Im}m_{H,\Lambda}(x+i\eta)=(\mathcal{P}_{\eta}f)(x)$.
@@
are Herglotz–Nevanlinna. Hence each $m_{\pm}$ is the Weyl coefficient of a (unique trace-normed) canonical system on $0,\infty)$. Moreover, the boundary densities satisfy $f^{\pm}(x)=\frac{1}{\pi}\operatorname{Im}m_{\pm}(x+i0)$ for a.e. $x\in\mathbb{R}$.
@@ 10.5. A quantitative minimax obstruction for de-smoothing
For $\eta>0$, recall the Poisson kernel (as in Corollary˜10.9)
@@ 11. An operator-theoretic corollary: compact perturbations of the free Dirac system
acting in $L^{2}_{H}$. This is the standard Dirac-type interpretation of canonical systems; the Weyl function $m_{H,\Lambda}$ is exactly the Weyl–Titchmarsh function of $\tau_{H}$ for the boundary condition at $s=0$ and the $L^{2}_{H}$ condition at $\infty$. (For background, see de Branges theory and modern accounts such as [9, 18].)
@@
$f(x):=\frac{1}{\pi}\operatorname{Im}m(x+i0),\qquad g_{\eta}(x):=\frac{1}{\pi}\operatorname{Im}m(x+i\eta).$
If in Theorem˜10.8 one has $c_{2}=0$ and $d\rho(t)=f(t)\,dt$, then combining Corollary˜10.9 with $P_{\eta}(x)=\frac{1}{\pi}\frac{\eta}{x^{2}+\eta^{2}}$ yields $g_{\eta}=\mathcal{P}_{\eta}f$. Then the map
@@
For the two-point alternatives $m^{\pm}$ produced by Theorem˜10.12 (realized via Theorem˜10.1), one has $c_{2}=0$ and $d\rho^{\pm}(t)=f^{\pm}(t)\,dt$, hence Corollary˜10.9 and the definition of $P_{\eta}$ give $g_{\eta}^{\pm}=\mathcal{P}_{\eta}f^{\pm}$. The two-point construction and deterministic minimax lower bound are exactly Theorem˜10.12. Finally, Theorem˜10.1 realizes the constructed Herglotz functions as Weyl functions of trace-normed canonical systems. ∎